# Quantum Fisher information matrices from Rényi relative entropies

### v1

@@ Abstract
Quantum generalizations of the Fisher information are important in quantum information science, with applications in high energy and condensed matter physics and in quantum estimation theory, machine learning, and optimization. One can derive a quantum generalization of the Fisher information matrix in a natural way as the Hessian matrix arising in a Taylor expansion of a smooth divergence. Such an approach is appealing and intuitive for quantum information theorists, given the ubiquity of divergences in quantum information theory. In contrast to the classical case, there is not a unique quantum generalization of the Fisher information matrix, similar to how there is not a unique quantum generalization of the relative entropy or the Rényi relative entropy. In this paper, I derive information matrices arising from the log-Euclidean, $\alpha$-$z$, and geometric Rényi relative entropies, with the main technical tool for doing so being the method of divided differences for calculating matrix derivatives. Interestingly, for all non-negative values of the Rényi parameter $\alpha$, the log-Euclidean Rényi relative entropy leads to the Kubo–Mori information matrix, and the geometric Rényi relative entropy leads to the right-logarithmic derivative Fisher information matrix. Thus, the resulting information matrices obey the data-processing inequality for all non-negative values of the Rényi parameter $\alpha$ even though the original quantities do not. Additionally, I derive and establish basic properties of $\alpha$-$z$ information matrices resulting from the $\alpha$-$z$ Rényi relative entropies, while leaving it open to determine all values of $\alpha$ and $z$ for which the $\alpha$-$z$ information matrices obey the data-processing inequality. For parameterized thermal states, I establish formulas for their $\alpha$-$z$ information matrices and hybrid quantum–classical algorithms for estimating them, with applications in quantum Boltzmann machine learning.
@@ 1.1 Background
One of the simplest ways for an information theorist to connect with the Fisher information concept is through smooth divergences that measure the distinguishability of nearby probability distributions. More precisely, the Fisher information matrix is equal to the Hessian matrix in a Taylor expansion of a smooth divergence evaluated at nearby probability distributions. Remarkably, the Čencov theorem states that, up to a constant prefactor, for every smooth divergence, the Fisher information matrix is the unique matrix that arises in this way [Čen82]. Thus, in the classical case, we can speak of a single notion of Fisher information matrix.
@@ 1.2 Summary of results
  • I establish ordering relations for the Petz– and sandwiched Rényi information matrices. Namely, as a function of the Rényi parameter $\alpha$, the Petz–Rényi information matrix is monotone decreasing on $\alpha\in\left(0,\frac{1}{2}\right]$ and monotone increasing on $\alpha\in\left[\frac{1}{2},\infty\right)$ (Theorem 27), and the sandwiched–Rényi information matrix is monotone increasing on $\alpha\in\left(0,\infty\right)$ (Theorem 29).
@@ 1.3 Paper organization
This paper is structured as follows. Section 2 reviews preliminary material needed to understand the rest of the paper, including classical Fisher information and quantum generalizations thereof. Section 3 presents the first result stated above, that the log-Euclidean information matrix is equal to the Kubo–Mori information matrix (Theorem 3). Section 4 presents the second result stated above, that the geometric information matrix is equal to the RLD Fisher information matrix (Theorem 5). Section 5 presents the formula for the $\alpha$-$z$ information matrix (Theorem 10), Section 6 presents special cases of the $\alpha$-$z$ information matrices, and Section 7 presents the ordering relations for the Petz– and sandwiched Rényi information matrices. Section 8 presents the formula for the $\alpha$-$z$ information matrix of parameterized thermal states (Theorem 31). Section 9 provides the decomposition formula for information matrices evaluated on classical-quantum states. Finally, Section 10 concludes with a summary of the results and points to directions for future work. Appendix B provides a detailed review of matrix derivatives.
@@ 2 Preliminaries
$\mathbb{R}_{++}\coloneqq\left(0,\infty\right).$ (2.1)
@@ 2.2 Quantum generalizations of Fisher information
Let $\left(\rho(\theta)\right)_{\theta\in\Theta}$ be a second-order differentiable parameterized family of positive definite states, where $\Theta\subseteq\mathbb{R}^{L}$ is open and $L\in\mathbb{N}$. For every smooth quantum divergence $\boldsymbol{D}$, there exists a function $\zeta\colon\mathbb{R}_{++}\times\mathbb{R}_{++}\to\mathbb{R}_{++}$ satisfying properties 1–4 below, such that the following equality holds:
$\left.\frac{\partial^{2}}{\partial\varepsilon_{i}\partial\varepsilon_{j}}\boldsymbol{D}(\rho(\theta)\|\rho(\theta+\varepsilon))\right|_{\varepsilon=0}=\left[I^{\zeta}(\theta)\right]_{i,j},$ (2.23)
where the quantum information matrix $I^{\zeta}(\theta)$ is defined in terms of its elements as
$\left[I^{\zeta}(\theta)\right]_{i,j}\coloneqq\sum_{k,\ell}\zeta(\lambda_{k},\lambda_{\ell})\operatorname{Tr}\!\left[\Pi_{k}\left(\partial_{i}\rho(\theta)\right)\Pi_{\ell}\left(\partial_{j}\rho(\theta)\right)\right],$ (2.24)
@@
    The function $t\mapsto f(t)\coloneqq\frac{1}{\zeta(t,1)}$ is operator monotone on the interval $t\in\left(0,\infty\right)$.
@@ 4.1 RLD Fisher information matrix from geometric Rényi relative entropies
Let $\left(\rho(\theta)\right)_{\theta\in\mathbb{R}^{L}}$ be a second-order differentiable parameterized family of positive definite states. The information matrix $\widehat{I}_{\alpha}(\theta)$ in (4.3) is equal to the RLD information matrix for all $\alpha\in\left(0,1\right)\cup\left(1,\infty\right)$:
@@ 4.2 RLD Fisher information matrix from Belavkin–Staszewski relative entropy
The Belavkin–Staszewski (Belavski) relative entropy is defined for positive definite states $\rho$ and $\sigma$ as follows [BS82]:
@@
The data-processing inequality holds for $\widehat{D}$ [HP91], and the Belavski relative entropy is known to be the following limit of the geometric Rényi relative entropy [Mat13, Mat18] (see also [KW21, Proposition 79]):
@@
Let us define the Belavski information matrix to be the information matrix induced by this relative entropy. That is, for a second-order differentiable parameterized family $\left(\rho(\theta)\right)_{\theta\in\mathbb{R}^{L}}$ of positive definite states, it is defined as
@@
Let $\left(\rho(\theta)\right)_{\theta\in\mathbb{R}^{L}}$ be a second-order differentiable parameterized family of positive definite states. The Belavski information matrix $\widehat{I}(\theta)$, defined in (4.66), is equal to the RLD information matrix:
@@ 5 $\alpha$-$z$ Information matrices from $\alpha$-$z$ Rényi relative entropies
Appendix D provides a long sequence of algebraic steps proving that
@@
After combining with (5.111) and incorporating the prefactor $\frac{z}{\alpha-1}$, the proof of (5.5) is concluded. For a proof of the case $x=y$ in (5.6), see Appendix C. ∎
@@ 6.2 Petz–Rényi information matrices
Consider that, for all $x,y>0$ such that $x\neq y$,
@@ 7 Orderings and relations between $\alpha$-$z$ information matrices
for all $\alpha\in\left(0,1\right)$. Additionally, it is known that, for fixed states $\rho$ and $\sigma$ and $\alpha>0$, the function $z\mapsto D_{\alpha,z}(\rho\|\sigma)$ is monotone increasing on $\mathbb{R}_{++}$ if $\alpha\in\left(0,1\right)$ and it is monotone decreasing on $\mathbb{R}_{++}$ if $\alpha>1$ [LT15, Proposition 1]. A direct consequence is the following:
@@
For all $\theta\in\mathbb{R}^{L}$, the matrix-valued function $z\mapsto I_{\alpha,z}(\theta)$ is monotone increasing on $\mathbb{R}_{++}$ if $\alpha\in\left(0,1\right)$ and it is monotone decreasing on $\mathbb{R}_{++}$ if $\alpha>1$, with respect to the Loewner order.
@@ 7.1 Ordering of general information matrices
Suppose that $\zeta\colon\mathbb{R}_{++}\times\mathbb{R}_{++}\to\mathbb{R}_{++}$ is a function satisfying items 1–3 of Theorem 1, and define $f(t)\coloneqq\frac{1}{\zeta(t,1)}$ for all $t>0$. Recall that a general information matrix has the following form:
$\left[I^{\zeta}(\theta)\right]_{i,j}$ $=\sum_{k,\ell}\zeta(\lambda_{k},\lambda_{\ell})\operatorname{Tr}\!\left[\Pi_{k}\left(\partial_{i}\rho(\theta)\right)\Pi_{\ell}\left(\partial_{j}\rho(\theta)\right)\right],$ (7.5)
@@
In this section, I drop the requirement for $f$ to be operator monotone on $\left(0,\infty\right)$, in which case $I^{\zeta}(\theta)$ need not satisfy the data-processing inequality. However, the information matrix $I^{\zeta}(\theta)$ is still positive semi-definite, as seen in Corollary 26 below, and we can still consider ordering relations between information matrices.
Let us begin with a general statement regarding ordering of information matrices built from functions $f_{1}$ and $f_{2}$ that satisfy an ordering relation on $\mathbb{R}_{++}$:
@@
Let $\zeta_{1},\zeta_{2}\colon\mathbb{R}_{++}\times\mathbb{R}_{++}\to\mathbb{R}_{++}$ be functions that satisfy items 1–3 of Theorem 1, and define $f_{i}(t)\coloneqq\frac{1}{\zeta_{i}(t,1)}$ for all $t>0$ and $i\in\left\{1,2\right\}$. Suppose that $f_{1}(t)\geq f_{2}(t)$ for all $t>0$. Then the following matrix inequality holds for all $\theta\in\mathbb{R}^{L}$:
$I^{\zeta_{1}}(\theta)\leq I^{\zeta_{2}}(\theta).$ (7.8)
@@
$v^{T}I^{\zeta_{1}}(\theta)v\leq v^{T}I^{\zeta_{2}}(\theta)v$ (7.9)
@@
$v^{T}I^{\zeta}(\theta)v$ $=\sum_{i,j}v_{i}\left(\sum_{k,\ell}\frac{1}{\lambda_{\ell}f\!\left(\frac{\lambda_{k}}{\lambda_{\ell}}\right)}\operatorname{Tr}\!\left[\Pi_{k}\left(\partial_{i}\rho(\theta)\right)\Pi_{\ell}\left(\partial_{j}\rho(\theta)\right)\right]\right)v_{j}$ (7.13)
@@
$v^{T}I^{\zeta_{1}}(\theta)v$ $=\langle W|\left[\sum_{k,\ell}\frac{1}{\lambda_{\ell}f_{1}\!\left(\frac{\lambda_{k}}{\lambda_{\ell}}\right)}\left(\Pi_{k}\otimes\Pi_{\ell}^{T}\right)\right]|W\rangle$ (7.23)
@@
Let $\zeta\colon\mathbb{R}_{++}\times\mathbb{R}_{++}\to\mathbb{R}_{++}$ be a function that satisfies items 1–3 of Theorem 1, and define $f(t)\coloneqq\frac{1}{\zeta(t,1)}$ for all $t>0$. Then the following matrix inequality holds for all $\theta\in\mathbb{R}^{L}$:
$I^{\zeta}(\theta)\geq 0.$ (7.27)
@@
$v^{T}I^{\zeta_{1}}(\theta)v$ $=\langle W|\left[\sum_{k,\ell}\frac{1}{\lambda_{\ell}f\!\left(\frac{\lambda_{k}}{\lambda_{\ell}}\right)}\left(\Pi_{k}\otimes\Pi_{\ell}^{T}\right)\right]|W\rangle.$ (7.28)

### v2

@@ Abstract
Quantum generalizations of the Fisher information are important in quantum information science, with applications in high energy and condensed matter physics and in quantum estimation theory, machine learning, and optimization. One can derive a quantum generalization of the Fisher information matrix in a natural way as the Hessian matrix arising in a Taylor expansion of a smooth divergence. Such an approach is appealing and intuitive for quantum information theorists, given the ubiquity of divergences in quantum information theory. In contrast to the classical case, there is not a unique quantum generalization of the Fisher information matrix, similar to how there is not a unique quantum generalization of the relative entropy or the Rényi relative entropy. In this paper, I derive information matrices arising from the log-Euclidean, $\alpha$-$z$, and geometric Rényi relative entropies, with the main technical tool for doing so being the method of divided differences for calculating matrix derivatives. Interestingly, for all non-negative values of the Rényi parameter $\alpha$, the log-Euclidean Rényi relative entropy leads to the Kubo–Mori information matrix, and the geometric Rényi relative entropy leads to the right-logarithmic derivative Fisher information matrix. Thus, the resulting information matrices obey the data-processing inequality for all non-negative values of the Rényi parameter $\alpha$ even though the original quantities do not. Additionally, I derive and establish basic properties of $\alpha$-$z$ information matrices resulting from the $\alpha$-$z$ Rényi relative entropies, while leaving it open to determine all values of $\alpha$ and $z$ for which the $\alpha$-$z$ information matrices obey the data-processing inequality. For parameterized thermal states and time-evolved states, I establish formulas for their $\alpha$-$z$ information matrices and hybrid quantum–classical algorithms for estimating them, with applications in quantum Boltzmann machine learning.
@@ 1.1 Background
One of the simplest ways for an information theorist to connect with the Fisher information concept is through smooth divergences that measure the distinguishability of nearby probability distributions. More precisely, the Fisher information matrix is equal to the Hessian matrix in a Taylor expansion of a smooth divergence evaluated at nearby probability distributions. Remarkably, the Čencov (Chentsov) theorem states that, up to a constant prefactor, for every smooth divergence, the Fisher information matrix is the unique matrix that arises in this way [Čen82]. Thus, in the classical case, we can speak of a single notion of Fisher information matrix.
@@ 1.2 Summary of results
  • I establish ordering relations for the Petz– and sandwiched Rényi information matrices. Namely, as a function of the Rényi parameter $\alpha$, the Petz–Rényi information matrix is monotone decreasing on $\alpha\in\left(0,\frac{1}{2}\right]$ and monotone increasing on $\alpha\in\left[\frac{1}{2},\infty\right)$ (Theorem 28), and the sandwiched Rényi information matrix is monotone increasing on $\alpha\in\left(0,\infty\right)$ (Theorem 30).
@@
  • For all $\alpha\in\left(0,1\right)$ and $z>0$, I derive a formula for the $\alpha$-$z$ information matrix when the underlying parameterized family consists of time-evolved states (Theorem 35). Specifically, for $\sigma(\phi)\coloneqq e^{-iH(\phi)}\rho e^{iH(\phi)}$, where $H(\phi)\coloneqq\sum_{j=1}^{L}\phi_{j}H_{j}$, with $\phi_{j}\in\mathbb{R}$ and $H_{j}$ Hermitian for all $j\in\left\{1,\ldots,L\right\}$, and $\rho\coloneqq\frac{e^{-G}}{\operatorname{Tr}\left[e^{-G}\right]}$ with $G$ Hermitian, the following formula holds for the $\alpha$-$z$ information matrix:
    $\left[I_{\alpha,z}(\phi)\right]_{i,j}=\left\langle\left[\Phi_{p_{\alpha,z}}\!\left(\Psi_{\phi}\!\left(H_{i}\right)\right),\left[G,\Psi_{\phi}(H_{j})\right]\right]\right\rangle_{\rho},$ (1.4)
    where the quantum channels $\Psi_{\phi}$ and $\Phi_{p_{\alpha,z}}$ are defined as
    $\Psi_{\phi}(X)$ $\coloneqq\int_{0}^{1}dt\ e^{iH(\phi)t}Xe^{-iH(\phi)t},$ (1.5)
    $\Phi_{p_{\alpha,z}}\!\left(X\right)$ $\coloneqq\int_{-\infty}^{\infty}dt\>p_{\alpha,z}(t)e^{-iGt}Xe^{iGt},$ (1.6)
    and $p_{\alpha,z}$ is the probability density function defined in (1.3).
@@
Similar to what was stated above for parameterized thermal states, the result in (1.4), regarding time-evolved states, is also a broad generalization of the findings reported in [MPW25a, Theorems 11, 15, and 19], such that these earlier findings are now special cases of Theorem 35. Specifically,
  • [MPW25a, Theorem 11] follows from Theorem 35 with $\alpha=z=\frac{1}{2}$.
  • [MPW25a, Theorem 19] follows from Theorem 35 with $\alpha\to 1$ and $z>0$ (or $\alpha>0$ and $z\to\infty$).
Given that time-evolved states are also known as quantum evolution machines [MPW25a], the result stated in (1.4) is also applicable to quantum machine learning. In particular, if one wishes to perform quantum optimization using natural gradient descent with quantum evolution machines and the geometry induced by the $\alpha$-$z$ Rényi relative entropy, then the formula in (1.4) is applicable. To estimate the elements of the $\alpha$-$z$ information matrix by means of a hybrid quantum-classical algorithm, one can employ a procedure similar to that outlined in [MPW25a, Figure 2(b)], with the underlying probability density $p$ therein replaced by $p_{\alpha,z}$, as defined in (1.3).
The results in (1.1) and (1.4) are also reminiscent of the classical Čencov theorem. Indeed, when considering the $\alpha$-$z$ information matrices of parameterized thermal states and time-evolved states for $\alpha\in\left(0,1\right)$ and $z>0$, the only $\alpha$-$z$ dependence that these expressions have is with respect to the underlying probability density, which is $q_{\alpha,z}$ in (1.1) and $p_{\alpha,z}$ in (1.4). Thus, these results represent a near unification of the $\alpha$-$z$ information matrices for these classes of states.
@@ 1.3 Paper organization
This paper is structured as follows. Section 2 reviews preliminary material needed to understand the rest of the paper, including classical Fisher information and quantum generalizations thereof. Section 3 presents the first result stated above, that the log-Euclidean information matrix is equal to the Kubo–Mori information matrix (Theorem 3). Section 4 presents the second result stated above, that the geometric information matrix is equal to the RLD Fisher information matrix (Theorem 5). Section 5 presents the formula for the $\alpha$-$z$ information matrix (Theorem 10), Section 6 presents special cases of the $\alpha$-$z$ information matrices, and Section 7 presents the ordering relations for the Petz– and sandwiched Rényi information matrices. Section 8 presents the formula for the $\alpha$-$z$ information matrix of parameterized thermal states (Theorem 32), and Section 9 presents the formula for the $\alpha$-$z$ information matrix of time-evolved states (Theorem 35). Section 10 provides the decomposition formula for information matrices evaluated on classical-quantum states. Finally, Section 11 concludes with a summary of the results and points to directions for future work. Appendix B provides a detailed review of matrix derivatives.
@@ 2 Preliminaries
$\mathbb{R}_{+}\coloneqq\left(0,\infty\right).$ (2.1)
@@ 2.2 Quantum generalizations of Fisher information
Let $\left(\rho(\theta)\right)_{\theta\in\Theta}$ be a second-order differentiable parameterized family of positive definite states, where $\Theta\subseteq\mathbb{R}^{L}$ is open and $L\in\mathbb{N}$. For every smooth quantum divergence $\boldsymbol{D}$, there exists a function $\zeta\colon\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ satisfying properties 1–4 below, such that the following equality holds:
$\left.\frac{\partial^{2}}{\partial\varepsilon_{i}\partial\varepsilon_{j}}\boldsymbol{D}(\rho(\theta)\|\rho(\theta+\varepsilon))\right|_{\varepsilon=0}=\left[I_{\zeta}(\theta)\right]_{i,j},$ (2.23)
where the quantum information matrix $I_{\zeta}(\theta)$ is defined in terms of its elements as
$\left[I_{\zeta}(\theta)\right]_{i,j}\coloneqq\sum_{k,\ell}\zeta(\lambda_{k},\lambda_{\ell})\operatorname{Tr}\!\left[\Pi_{k}\left(\partial_{i}\rho(\theta)\right)\Pi_{\ell}\left(\partial_{j}\rho(\theta)\right)\right],$ (2.24)
@@
    The function $t\mapsto f(t)\coloneqq\frac{1}{\zeta(t,1)}$ is operator monotone on $t\in\left(0,\infty\right)$.
@@ 4.1 RLD Fisher information matrix from geometric Rényi relative entropies
Let $\left(\rho(\theta)\right)_{\theta\in\mathbb{R}^{L}}$ be a second-order differentiable parameterized family of positive definite states. The geometric information matrix $\widehat{I}_{\alpha}(\theta)$ in (4.3) is equal to the RLD information matrix for all $\alpha\in\left(0,1\right)\cup\left(1,\infty\right)$:
@@ 4.2 RLD Fisher information matrix from Belavkin–Staszewski relative entropy
The Belavkin–Staszewski relative entropy is defined for positive definite states $\rho$ and $\sigma$ as follows [BS82]:
@@
The data-processing inequality holds for $\widehat{D}$ [HP91], and the Belavkin–Staszewski relative entropy is known to be the following limit of the geometric Rényi relative entropy [Mat13, Mat18] (see also [KW21, Proposition 79]):
@@
Let us define the Belavkin–Staszewski information matrix to be the information matrix induced by this relative entropy. That is, for a second-order differentiable parameterized family $\left(\rho(\theta)\right)_{\theta\in\mathbb{R}^{L}}$ of positive definite states, it is defined as
@@
Let $\left(\rho(\theta)\right)_{\theta\in\mathbb{R}^{L}}$ be a second-order differentiable parameterized family of positive definite states. The Belavkin–Staszewski information matrix $\widehat{I}(\theta)$, defined in (4.66), is equal to the RLD information matrix:
@@ 5 $\alpha$-$z$ Information matrices from $\alpha$-$z$ Rényi relative entropies
Appendix C provides a long sequence of algebraic steps proving that
@@
After combining with (5.111) and incorporating the prefactor $\frac{z}{\alpha-1}$, the proof of (5.5) is concluded. For a proof of the case $x=y$ in (5.6), see Appendix D. ∎
@@
Remark 13.
For two pure states $\psi\equiv|\psi\rangle\!\langle\psi|$ and $\phi\equiv|\phi\rangle\!\langle\phi|$ and $\alpha\in\left(0,1\right)$, the $\alpha$-$z$ Rényi relative entropy reduces as follows:
$D_{\alpha,z}(\psi\|\phi)$ $\coloneqq\frac{1}{\alpha-1}\ln\operatorname{Tr}\!\left[\left(\phi^{\frac{1-\alpha}{2z}}\psi^{\frac{\alpha}{z}}\phi^{\frac{1-\alpha}{2z}}\right)^{z}\right]$ (5.137)
$=\frac{1}{\alpha-1}\ln\operatorname{Tr}\!\left[\left(|\phi\rangle\langle\phi|\psi\rangle\langle\psi|\phi\rangle\langle\phi|\right)^{z}\right]$ (5.138)
$=\frac{1}{\alpha-1}\ln\operatorname{Tr}\!\left[\left(|\phi\rangle\left|\langle\phi|\psi\rangle\right|^{2}\langle\phi|\right)^{z}\right]$ (5.139)
$=\frac{1}{\alpha-1}\ln\left|\langle\phi|\psi\rangle\right|^{2z}\operatorname{Tr}\!\left[\left(|\phi\rangle\!\langle\phi|\right)^{z}\right]$ (5.140)
$=\frac{1}{\alpha-1}\ln\left|\langle\phi|\psi\rangle\right|^{2z}\operatorname{Tr}\!\left[|\phi\rangle\!\langle\phi|\right]$ (5.141)
$=\frac{1}{\alpha-1}\ln\left|\langle\phi|\psi\rangle\right|^{2z}$ (5.142)
$=\frac{z}{1-\alpha}\left(-\ln\left|\langle\phi|\psi\rangle\right|^{2}\right)$ (5.143)
$=\frac{z}{1-\alpha}\left(-\ln F(\psi,\phi)\right),$ (5.144)
where $F(\psi,\phi)\coloneqq\left|\langle\phi|\psi\rangle\right|^{2}$ is the fidelity. As expected, for $\psi\neq\phi$, the limit $D_{\alpha,z}(\psi\|\phi)\to\infty$ holds as $\alpha\to 1$ for fixed $z>0$ or $D_{\alpha,z}(\psi\|\phi)\to\infty$ as $z\to\infty$ for fixed $\alpha\in\left(0,1\right)$, consistent with the fact that $D_{\alpha,z}$ converges to the standard (Umegaki) relative entropy in these limits and this latter quantity being equal to $\infty$ for two distinct pure states. Thus, for a parameterized family $\left(\psi(\theta)\right)_{\theta\in\mathbb{R}^{L}}$ of pure states, the $\alpha$-$z$ information matrix reduces to
$\left[I_{\alpha,z}(\theta)\right]_{i,j}$ $\coloneqq\frac{1}{\alpha}\left.\frac{\partial^{2}}{\partial\varepsilon_{i}\partial\varepsilon_{j}}D_{\alpha,z}(\psi(\theta)\|\psi(\theta+\varepsilon))\right|_{\varepsilon=0}.$ (5.145)
$=\frac{z}{\alpha\left(1-\alpha\right)}\left.\frac{\partial^{2}}{\partial\varepsilon_{i}\partial\varepsilon_{j}}\left(-\ln F(\psi(\theta),\psi(\theta+\varepsilon))\right)\right|_{\varepsilon=0}$ (5.146)
$=\frac{2z}{\alpha\left(1-\alpha\right)}\Re\left[\langle\partial_{i}\psi(\theta)|\left(I-|\psi(\theta)\rangle\!\langle\psi(\theta)|\right)|\partial_{j}\psi(\theta)\rangle\right],$ (5.147)
consistent with the known result [LYLW19, Theorem 2.5] that
$\left.\frac{\partial^{2}}{\partial\varepsilon_{i}\partial\varepsilon_{j}}\left(-2\ln F(\psi(\theta),\psi(\theta+\varepsilon))\right)\right|_{\varepsilon=0}=4\Re\left[\langle\partial_{i}\psi(\theta)|\left(I-|\psi(\theta)\rangle\!\langle\psi(\theta)|\right)|\partial_{j}\psi(\theta)\rangle\right].$ (5.148)
Appendix E reviews the proof of (5.148).
@@ 6.2 Petz–Rényi information matrices
Considering that, for all $x,y>0$ such that $x\neq y$,
@@ 7 Orderings and relations between $\alpha$-$z$ information matrices
for all $\alpha\in\left(0,1\right)$. Additionally, it is known that, for fixed states $\rho$ and $\sigma$ and $\alpha>0$, the function $z\mapsto D_{\alpha,z}(\rho\|\sigma)$ is monotone increasing on $\mathbb{R}_{+}$ if $\alpha\in\left(0,1\right)$ and it is monotone decreasing on $\mathbb{R}_{+}$ if $\alpha>1$ [LT15, Proposition 1]. A direct consequence is the following:
@@
For all $\theta\in\mathbb{R}^{L}$, the matrix-valued function $z\mapsto I_{\alpha,z}(\theta)$ is monotone increasing on $\mathbb{R}_{+}$ if $\alpha\in\left(0,1\right)$ and it is monotone decreasing on $\mathbb{R}_{+}$ if $\alpha>1$, with respect to the Loewner order.
@@ 7.1 Ordering of general information matrices
Suppose that $\zeta\colon\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ is a function satisfying items 1–3 of Theorem 1, and define $f(t)\coloneqq\frac{1}{\zeta(t,1)}$ for all $t>0$. Recall that a general information matrix has the following form:
$\left[I_{\zeta}(\theta)\right]_{i,j}$ $=\sum_{k,\ell}\zeta(\lambda_{k},\lambda_{\ell})\operatorname{Tr}\!\left[\Pi_{k}\left(\partial_{i}\rho(\theta)\right)\Pi_{\ell}\left(\partial_{j}\rho(\theta)\right)\right],$ (7.5)
@@
In this section, I drop the requirement for $f$ to be operator monotone on $\left(0,\infty\right)$, in which case $I_{\zeta}(\theta)$ need not satisfy the data-processing inequality. However, the information matrix $I_{\zeta}(\theta)$ is still positive semi-definite, as seen in Corollary 27 below, and we can still consider ordering relations between information matrices.
Let us begin with a general statement regarding ordering of information matrices built from functions $f_{1}$ and $f_{2}$ that satisfy an ordering relation on $\mathbb{R}_{+}$:
@@
Let $\zeta_{1},\zeta_{2}\colon\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ be functions that satisfy items 1–3 of Theorem 1, and define $f_{i}(t)\coloneqq\frac{1}{\zeta_{i}(t,1)}$ for all $t>0$ and $i\in\left\{1,2\right\}$. Suppose that $f_{1}(t)\geq f_{2}(t)$ for all $t>0$. Then the following matrix inequality holds for all $\theta\in\mathbb{R}^{L}$:
$I_{\zeta_{1}}(\theta)\leq I_{\zeta_{2}}(\theta).$ (7.8)
@@
$v^{T}I_{\zeta_{1}}(\theta)v\leq v^{T}I_{\zeta_{2}}(\theta)v$ (7.9)
@@
$v^{T}I_{\zeta}(\theta)v$ $=\sum_{i,j}v_{i}\left(\sum_{k,\ell}\frac{1}{\lambda_{\ell}f\!\left(\frac{\lambda_{k}}{\lambda_{\ell}}\right)}\operatorname{Tr}\!\left[\Pi_{k}\left(\partial_{i}\rho(\theta)\right)\Pi_{\ell}\left(\partial_{j}\rho(\theta)\right)\right]\right)v_{j}$ (7.13)
@@
$v^{T}I_{\zeta_{1}}(\theta)v$ $=\langle W|\left[\sum_{k,\ell}\frac{1}{\lambda_{\ell}f_{1}\!\left(\frac{\lambda_{k}}{\lambda_{\ell}}\right)}\left(\Pi_{k}\otimes\Pi_{\ell}^{T}\right)\right]|W\rangle$ (7.23)
@@
Let $\zeta\colon\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a function that satisfies items 1–3 of Theorem 1, and define $f(t)\coloneqq\frac{1}{\zeta(t,1)}$ for all $t>0$. Then the following matrix inequality holds for all $\theta\in\mathbb{R}^{L}$:
$I_{\zeta}(\theta)\geq 0.$ (7.27)
@@
$v^{T}I_{\zeta_{1}}(\theta)v$ $=\langle W|\left[\sum_{k,\ell}\frac{1}{\lambda_{\ell}f\!\left(\frac{\lambda_{k}}{\lambda_{\ell}}\right)}\left(\Pi_{k}\otimes\Pi_{\ell}^{T}\right)\right]|W\rangle.$ (7.28)